# On homomorphic images of ultraproducts

### v1

@@ 1. Introduction
If $\mathcal{U}$ is a nonprinciple ultrafilter on $\omega$ and $H$ is a group such that every $h\in H$ lies in a homomorphic image within $H$ of $(\prod_{\omega}\mathbb{Z})/\mathcal{U}$, must $H$ be a homomorphic image of a nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{V}$?
@@
Questions 1 and 2 will be answered affirmatively using some new characterizations of cotorsion abelian groups (Theorem 16). We’ll see that the direct sum $\bigoplus_{n\in\omega\setminus\{0\}}\mathbb{Z}/n\mathbb{Z}$ examplifies a negative answer to Question 3, and many more examples will follow from Corollary 22. It was added in the proof of the article [1] that Question 4 has a positive solution; we provide a very short argument using the machinery presented. It may not be immediately clear what Questions 2 and 4 have to do with ultraproducts, but it will be seen that all these questions are of a kind. The principal tool in attacking Questions 1 - 4 is the notion of Higman-completeness, due to Herefort and Hojka [10]. We will not directly answer Question 5, but will show that an affirmative answer is consistent with ZFC. Indeed, a positive answer follows immediately from Blass’s near coherence of filters [2].
@@ 2. Higman-complete groups
We begin this section by defining a handful of concepts and state a proposition detailing some known relationships among them. Next, we pass through some lemmas towards Theorem 16, answering some questions along the way. We emphasize that although much of the discourse will involve abelian groups, we do not require any of our groups to be abelian unless otherwise stated.
@@
Assume the hypotheses. We will show that $A$ is Higman-complete, and the fact that $A$ is cotorsion then follows immediately from Proposition 10. Let $(f_{i})_{i\in\omega}$ be a sequence of elements of $A$ and $(w_{i})_{i\in\omega}$ a sequence of words in two variables. Since $A$ is abelian, we can without loss of generality assume that the word $w_{i}(x,y)$ is of form $z_{i}x+z_{i}^{\prime}y$ where $z_{i},z_{i}^{\prime}\in\mathbb{Z}$. For each $i\in\omega$, let $\delta_{i}\in\mathbb{Z}^{\omega}$ be the element with $\delta_{i}(i)=1$ and $\delta_{i}(j)=0$ for $j\neq i$. Letting $\phi:\bigoplus_{\omega}\mathbb{Z}\rightarrow A$ be defined by the function $\delta_{i}\mapsto f_{i}$, we obtain (by assumption) an extension $\Phi:\mathbb{Z}^{\omega}\rightarrow A$. Letting
@@
it is clear that the sequence $(h_{i})_{i\in\omega}$ satisfies $h_{i}=w_{i}(f_{i},h_{i+1})$. Thus $A$ is Higman-complete, and therefore cotorsion by Proposition 10.
@@
Let a sequence $([\zeta_{i}])_{i\in\omega}$ of elements of $G$ be given, together with a sequence $(w_{i})_{i\in\omega}$ of free words in two variables. For each $i\in\omega$ we fix a representative $\zeta_{i}\in[\zeta_{i}]$. We shall define by induction a sequence $(\eta_{i})_{i\in\omega}$. Let $\eta_{0}(0)=w_{0}(\zeta_{0}(0),1_{G_{0}})$. Let $\eta_{1}(1)=w_{1}(\zeta_{1}(1),1_{G_{1}})$ and $\eta_{0}(1)=w_{0}(\zeta_{0}(1),\eta_{1}(1))$. Generally, if $\eta_{i}(j)$ has been defined for all $i\leq j\leq k$, then we let
@@
For $j<i$ we let $\eta_{i}(j)=1_{G_{j}}$.
We claim that the sequence $([\eta_{i}])_{i\in\omega}$ of elements in $G$ verifies the definition of Higman-completeness. To see this, let $i\in\omega$ be given. By construction we have for all $j\geq i$ that $\eta_{i}(j)=w_{i}(\zeta_{i}(j),\eta_{i+1}(j))$, so that we can indeed write $[\eta_{i}]=w_{i}([\zeta_{i}],[\eta_{i+1}])$.
For the claim in the second sentence of this lemma we note that $(\prod_{i\in\omega}G_{i})/\mathcal{U}$ is a quotient of $(\prod_{i\in\omega}G_{i})/\mathcal{U}$ and apply Remark 9.
Proof of negative solution to Question 3.
@@
$H=\bigoplus_{n\in\omega\setminus\{0\}}\mathbb{Z}/n\mathbb{Z}.$
This group is torsion, and so each $h\in H$ is an element of a finite cyclic subgroup $C$ of $H$, and it a trivial matter to surject $(\prod_{\omega}\mathbb{Z})/\mathcal{U}$ onto $C$. Suppose for contradiction that $H$ is a homomorphic image of a nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{V}$. Then $H$ is Higman-complete (by Lemma 11 and Remark 9). Thus $H$ is an abelian group which is Higman-complete, and so $H$ is cotorsion (Proposition 10). Since $H$ is a torsion cotorsion group, it is a direct sum of a bounded group (i.e. a group in which $a^{n}=0$ for some universal $n\in\omega$) and a divisible group ([8, Corollary 9.8.4]). However $H$ has no nontrivial divisible elements and is not bounded, a contradiction. We shall later see that many nonabelian groups of this flavor also serve as counterexamples. ∎
@@
    It is not the case that $G$ is perfect and of finite commutator length.
@@
(b) $\Rightarrow$ (a). Take $\mathcal{U}$ to be a nonprincipal ultrafilter on $\omega$. Let $H$ be the ultrapower $(\prod_{\omega}G)/\mathcal{U}$. Then $H$ is Higman-complete (Lemma 11). To see that the abelianization of $H$ is not trivial, select an element $(g_{i})_{i\in\omega}$ in $\prod_{i\in\omega}G$ such that the commutator length of $g_{i}$ is greater than $i$ (possibly $\infty$). Then the element $[(g_{i})_{i\in\omega}]$ in $H$ is not in the kernel of the abelianization of $H$ (since its commutator length is infinite). Therefore the abelianization of $H$ is nontrivial, and is Higman-complete as a homomorphic image of $H$, and therefore cotorsion. That $H$ is elementarily equivalent to $G$ is standard.
(a) $\Rightarrow$ (b) We prove the contrapositive. Supposing that $G$ is perfect and of bounded commutator length, say length $j$, every group $H$ which is elementarily equivalent to $G$ also is perfect of bounded commutator length at most $j$. Then the abelianization of such an $H$ is trivial. ∎
Continuing our progess toward Theorem 16 we give two more lemmas.
@@
Given a sequence $(\zeta_{i})_{i}$ of elements in $G$, and a sequence $(w_{i})_{i\in\omega}$ of free words in two variables, we solve the nececcary equations coordinatewise. More precisely, for a fixed $j\in J$ we have a sequence $(\zeta_{i}(j))_{i\in\omega}$ of elements of $G_{j}$ and we select $(\eta_{i}(j))_{i\in\omega}$ so that $\eta_{i}(j)=w_{i}(\zeta_{i}(j),\eta_{i+1}(j))$ for each $i\in\omega$ (which is possible since $G_{j}$ is Higman-complete). Now, it is clear that $\eta_{i}=w_{i}(\zeta_{i},\eta_{i+1})$ holds, since the equation holds in each coordinate. ∎
@@
The following lemma is probably known.
@@
Letting $D$ denote the divisible subgroup of $\mathbb{A}$, we can write $\mathbb{A}=D\oplus R$ where $R$ is a reduced algebraically compact group. A reduced algebraically compact group is a direct summand of a product of cyclic groups of prime power order [8, Corollary 6.1.4], so more particularly $R$ is a homomorphic image of a product of finite cyclic groups.
@@
Let $\kappa\geq|D|$ be an infinite cardinal. Then $\mathbb{Q}^{\kappa}\simeq\bigoplus_{i\in 2^{\kappa}}\mathbb{Q}_{i}$, where each $\mathbb{Q}_{i}$ is isomorphic to $\mathbb{Q}$. Let $f:2^{\kappa}\rightarrow D$ be a surjective function. For each $i\in 2^{\kappa}$ define a homomorphism $\phi_{i}:\mathbb{Q}_{i}\rightarrow D$ defined by having $\phi_{i}(1)=f(i)=:a_{1}$, $\phi_{i}(\frac{1}{2!})=a_{2}$ be an element of $D$ with $a_{2}^{2}=a_{1}$, $\phi_{i}(\frac{1}{3!})=a_{3}$ where $a_{3}^{3}=a_{2}$, $\phi_{i}(\frac{1}{4!})=a_{4}$ where $a_{4}^{4}=a_{3}$, etc. It is clear that this is indeed a homomorphism. Now let $\phi:\bigoplus_{i\in 2^{\kappa}}\mathbb{Q}_{i}\rightarrow D$ be the homomorphism defined by the $\phi_{i}$. By construction, $\phi$ is surjective, and the domain of $\phi$ is isomorphic to a product of copies of $\mathbb{Q}$. Thus $D$ is a homomorphic image of a product of finite cyclic groups, and as $R$ also satisfies this property, the group $\mathbb{A}=D\oplus R$ also has this property.
@@
(1) $\Rightarrow$ (2) Suppose that $A$ is cotorsion. By Proposition 10 we know that there is a sequence $(A_{i})_{i\in\omega}$ of abelian groups and a nonprincipal ultrafilter $\mathcal{U}$ on $\omega$ so that $A$ is a homomorphic image of the ultraproduct $(\prod_{i\in\omega}A_{i})/\mathcal{U}$. As $(\prod_{i\in\omega}A_{i})/\mathcal{U}$ is a homomorphic image of $(\prod_{i\in\omega}A_{i})/(\bigoplus_{i\in\omega}A_{i})$, so is $A$.
(1) $\Rightarrow$ (3) This follows from the fact that a cotorsion group is a homomorphic image of an algebraically compact group (Proposition 10) and the fact that an algebraically compact abelian group is a homomorphic image of a product of finite cyclic groups (Lemma 15).
(1) $\Rightarrow$ (4) This follows because (1) $\Rightarrow$ (3) $\Rightarrow$ (4).
(2) $\Rightarrow$ (1) If $A$ is a homomorphic image of a group of form $(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$ then $A$ is Higman-complete (Lemma 11), so $A$ is cotorsion (Proposition 10).
(4) $\Rightarrow$ (1) Note first that a finite group $F$ is Higman-complete. To see this, take a nonprincipal ultrafilter $\mathcal{V}$ on $\omega$ and recall that $F$ is isomorphic to the ultrapower $(\prod_{\omega}F)/\mathcal{V}$, so $F$ is Higman-complete (Proposition 11). Therefore a product of finite groups is Higman-complete (Lemma 13), and so $A$ is Higman-complete as the homomorphic image of a Higman-complete group (Remark 9). Thus $A$ is cotorsion (Proposition 10).
(3) $\Rightarrow$ (1) This follows because (3) $\Rightarrow$ (4) $\Rightarrow$ (1).
@@ 3. An analogue of a theorem of Chase
In light of the very abelian nature of many of the previous results, we give a general method for producing (nonabelian) examples which give a negative solution to Question 3. These will be torsion groups which are not the homomorphic image of any nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{U}$. For this we will prove an analogue of a result of Chase [7].
@@
We’ll say a group $H$ is bounded torsion if there exists some $n\in\omega$ such that $(\forall h\in H)h^{n}=1_{H}$. For such an $H$ we let $\operatorname{ex}(H)\in\omega$ denote the smallest such $n$. If a torsion group is not bounded then we shall say it is unbounded torsion.
@@
Let $\phi:G\rightarrow\textbf{Bdd}\{H_{j}\}_{j\in J}$ be an abstract group homomorphism, with $G$ a topological group which is either
@@
and each $H_{j}$ is a group which is bounded torsion. Then there is some open neighborhood $U$ of the identity $1_{G}$ and $N\in\omega$ such that $(\forall h\in\phi(U))h^{N}=1$.
@@
Let $H=\textbf{Bdd}\{H_{j}\}_{j\in J}$. Suppose for contradiction that the conclusion fails. The strategy will be to produce a system of equations (like in the definition of Higman-completeness) whose solution will give a contradiction. Suppose we are in situation (1) and let $d$ be a complete metric which induces the topology on $G$. We inductively choose
@@
For convenience, in the remainder of this section we’ll say a group which is a homomorphic image of a nonprincipal ultraproduct over $\omega$ is a ()(*)( ∗ )-group.
@@
If $\{H_{j}\}_{j\in J}$ is a collection of bounded torsion groups for which $\{\operatorname{ex}(H_{j})\}_{j\in J}$ is unbounded in $\omega$, then the direct sum $\bigoplus_{j\in J}H_{j}$ is not a ()(*)( ∗ )-group.
Part of the attraction of Corollary 22 is that it provides many examples of non-()(*)( ∗ )-groups which are (unbounded) torsion. One can apply Corollary 22 to the collection $\{S_{j}\}_{j\in\omega}$ where $S_{j}$ is the symmetric group on $j$ elements, or to the collection $\{\mathfrak{b}(2,j)\}_{j\in\omega\setminus\{0,1\}}$ where $\mathfrak{b}(2,j)$ is the free Burnside group of rank $2$ and exponent $j$.
There exist ()(*)( ∗ )-groups which are unbounded torsion. For example, a quasicyclic group $\mathbb{Z}(p^{\infty})$ is cotorsion and therefore a ()(*)( ∗ )-group. We do not know the answers to the following.
@@
    If a ()(*)( ∗ )-group is unbounded torsion then does it include a nontrivial divisible subgroup?
@@
    Is there a bounded torsion group which is not a ()(*)( ∗ )-group?
@@
We give a few more examples of non-()(*)( ∗ )-groups. Recall that a group $H$ is cm-slender if every abstract group homomorphism from a completely metrizable group to $H$ has open kernel [6]. A cm-slender group is necessarily torsion-free, and examples include Baumslag-Solitar groups and nontrivial word-hyperbolic groups. Arguing as in Corollary 21 it is clear that a group having nontrivial homomorphic image in a cm-slender group is not a ()(*)( ∗ )-group. Reasoning similarly, an infinite subgroup of the mapping class group of a connected compact surface is not a ()(*)( ∗ )-group [4, Theorem 9.1], nor is a group which is nontrivially a free product of groups [13, Theorem 1.5].

### v2

@@ 1. Introduction
If $\mathcal{U}$ is a nonprincipal ultrafilter on $\omega$ and $H$ is a group such that every $h\in H$ lies in a homomorphic image within $H$ of $(\prod_{\omega}\mathbb{Z})/\mathcal{U}$, must $H$ be a homomorphic image of a nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{V}$?
@@
Questions 1 and 2 will be answered affirmatively using some new characterizations of cotorsion abelian groups (Theorem 17). We’ll see that the direct sum $\bigoplus_{n\in\omega\setminus\{0\}}\mathbb{Z}/n\mathbb{Z}$ gives a negative answer to Question 3, and many more examples will follow from Corollary 23. It was added in the proof of the article [1] that Question 4 has a positive solution; we provide a very short argument using the machinery presented. It may not be immediately clear what Questions 2 and 4 have to do with ultraproducts, but it will be seen that all these questions are of a kind. The principal tool in attacking Questions 1 - 4 is the notion of Higman-completeness, due to Herefort and Hojka [10]. We will not completely answer Question 5, but will show that an affirmative answer is consistent with ZFC. Indeed, a positive answer follows from Blass’s near coherence of filters [2].
@@ 2. Higman-complete groups
We begin this section by defining a handful of concepts and state a proposition detailing some known relationships among them. Next, we pass through some lemmas towards Theorem 17, answering some questions along the way. We emphasize that although much of the discourse will involve abelian groups, we do not require any of our groups to be abelian unless so stated.
@@
Assume the hypotheses. We will show that $A$ is Higman-complete, and the fact that $A$ is cotorsion then follows immediately from Proposition 10. Let $(f_{i})_{i\in\omega}$ be a sequence of elements of $A$ and $(w_{i})_{i\in\omega}$ a sequence of words in two variables. Since $A$ is abelian, we can without loss of generality assume that the word $w_{i}(x,y)$ is of form $z_{i}x+z_{i}^{\prime}y$ where $z_{i},z_{i}^{\prime}\in\mathbb{Z}$. For each $i\in\omega$, let $\delta_{i}\in\mathbb{Z}^{\omega}$ be the element with $\delta_{i}(i)=1$ and $\delta_{i}(j)=0$ for $j\neq i$. Letting $\phi:\bigoplus_{\omega}\mathbb{Z}\rightarrow A$ be defined by the function $\delta_{i}\mapsto f_{i}$, we obtain (by the hypothesis of Question 4) an extension $\Phi:\mathbb{Z}^{\omega}\rightarrow A$. Letting
@@
it is easy to check that the sequence $(h_{i})_{i\in\omega}$ satisfies $h_{i}=w_{i}(f_{i},h_{i+1})$. For example,
$h_{0}=\Phi(z_{0}\delta_{0})+z_{0}^{\prime}\Phi(0,z_{1},z_{1}^{\prime}z_{2},\ldots)=z_{0}f_{0}+z_{0}^{\prime}h_{1}$
and for $i>0$ the check is entirely analogous. Thus $A$ is Higman-complete, and therefore cotorsion by Proposition 10.
@@
In this lemma we will use symbols $\zeta$ and $\eta$ to denote elements of $\prod_{i\in\omega}G_{i}$ and $[\zeta],[\eta]$ to denote elements of $G$. Let a sequence $([\zeta_{\ell}])_{\ell\in\omega}$ of elements of $G$ be given, together with a sequence $(w_{\ell})_{\ell\in\omega}$ of free words in two variables. For each $\ell\in\omega$ we fix a representative $\zeta_{\ell}\in[\zeta_{\ell}]$. We shall define by induction a sequence $(\eta_{\ell})_{\ell\in\omega}$ of elements in $\prod_{i\in\omega}G_{i}$. Let $\eta_{0}(0)=w_{0}(\zeta_{0}(0),1_{G_{0}})$. Let $\eta_{1}(1)=w_{1}(\zeta_{1}(1),1_{G_{1}})$ and $\eta_{0}(1)=w_{0}(\zeta_{0}(1),\eta_{1}(1))$. Generally, if $\eta_{\ell}(j)$ has been defined for all $\ell\leq j\leq k$, then we let
@@
For $j<\ell$ we let $\eta_{\ell}(j)=1_{G_{j}}$.
We claim that the sequence $([\eta_{\ell}])_{\ell\in\omega}$ of elements in $G$ satisfies the definition of Higman-completeness. To see this, let $\ell\in\omega$ be given. By construction we have for all $j\geq\ell$ that $\eta_{\ell}(j)=w_{\ell}(\zeta_{\ell}(j),\eta_{\ell+1}(j))$, so that we can indeed write $[\eta_{\ell}]=w_{\ell}([\zeta_{\ell}],[\eta_{\ell+1}])$.
For the claim in the second sentence of this lemma we note that $(\prod_{i\in\omega}G_{i})/\mathcal{U}$ is a quotient of $(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$ and apply Remark 9.
Proof of negative answer to Question 3.
@@
$H=\bigoplus_{n\in\omega}C_{n}$
where the $C_{n}$ are finite cyclic groups of unbounded orders. This group is torsion, and so each $h\in H$ is an element of a finite cyclic subgroup $C$ of $H$, say $|C|=m$, and it a trivial matter to surject $(\prod_{\omega}\mathbb{Z})/\mathcal{U}$ onto $(\prod_{\omega}\mathbb{Z}/m\mathbb{Z})/\mathcal{U}$ and as $m$ is finite we have $(\prod_{\omega}\mathbb{Z}/m\mathbb{Z})/\mathcal{U}$ isomorphic to $C$. Suppose by way of contradiction that $H$ is a homomorphic image of a nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{V}$. Then $H$ is Higman-complete (by Lemma 11 and Remark 9). Thus $H$ is an abelian group which is Higman-complete, and so $H$ is cotorsion (Proposition 10). Since $H$ is a torsion cotorsion group, it is a direct sum of a bounded group (i.e. a group in which $a^{n}=0$ for some universal $n\in\omega$) and a divisible group ([8, Corollary 9.8.4]). However $H$ has no nontrivial divisible elements and is not bounded, a contradiction. We shall later see that many nonabelian groups of this flavor also serve as counterexamples. ∎
@@
    Either $G$ is not perfect or $G$ has elements of unbounded commutator length.
@@
(b) $\Rightarrow$ (a). Take $\mathcal{U}$ to be a nonprincipal ultrafilter on $\omega$. Let $H$ be the ultrapower $(\prod_{\omega}G)/\mathcal{U}$. Then $H$ is Higman-complete (Lemma 11), elementarily equivalent to $G$, and the abelianization $H/H^{\prime}$ is cotorsion. It remains to prove that $H/H^{\prime}$ is nontrivial. If $G$ is not perfect, then the natural surjective map from $H=(\prod_{\omega}G)/\mathcal{U}$ to $(\prod_{\omega}G/G^{\prime})/\mathcal{U}$ gives a nontrivial abelian image of $H$, so $H/H^{\prime}$ is nontrivial. On the other hand if $G$ is perfect then $G$ has elements of arbitrarily long commutator length, so select an element $(g_{i})_{i\in\omega}$ in $\prod_{i\in\omega}G$ such that the commutator length of $g_{i}$ is greater than $i$. Then the element $[(g_{i})_{i\in\omega}]$ in $H$ is not in the kernel of the abelianization of $H$ (since its commutator length is infinite).
(a) $\Rightarrow$ (b) We prove the contrapositive. Supposing that $G$ is perfect and of bounded commutator length, say length $j$, every group $H$ which is elementarily equivalent to $G$ also is perfect of bounded commutator length at most $j$ (since this property can be expressed using a first-order formula). Then the abelianization of such an $H$ is trivial. ∎
Continuing our progress toward Theorem 17 we give two more lemmas.
@@
Given a sequence $(\zeta_{i})_{i}$ of elements in $G$, and a sequence $(w_{i})_{i\in\omega}$ of free words in two variables, we solve the necessary equations coordinatewise. More precisely, for a fixed $j\in J$ we have a sequence $(\zeta_{i}(j))_{i\in\omega}$ of elements of $G_{j}$ and we select $(\eta_{i}(j))_{i\in\omega}$ so that $\eta_{i}(j)=w_{i}(\zeta_{i}(j),\eta_{i+1}(j))$ for each $i\in\omega$ (which is possible since $G_{j}$ is Higman-complete). Now, it is clear that $\eta_{i}=w_{i}(\zeta_{i},\eta_{i+1})$ holds, since the equation holds in each coordinate. ∎
@@
Lemma 15 below is immediate from [8, Theorem 4.3.1], and Lemma 16 is probably known.
Lemma 15.
An abelian group $D$ is divisible if and only if $D$ is the homomorphic image of $\bigoplus_{\kappa}\mathbb{Q}$ for some cardinal $\kappa$.
@@
Letting $D$ denote the maximal divisible subgroup of $\mathbb{A}$, we can write $\mathbb{A}=D\oplus R$ where $R$ is a reduced algebraically compact group. A reduced algebraically compact group is a direct summand of a product of cyclic groups of prime power order [8, Corollary 6.1.4], so in particular $R$ is a homomorphic image of a product of finite cyclic groups.
@@
By Lemma 15 take $\kappa$ to be a cardinal such that $D$ is a homomorphic image of $\bigoplus_{\kappa}\mathbb{Q}$. Now $\bigoplus_{\kappa}\mathbb{Q}$ is a homomorphic image of $\mathbb{Q}^{\kappa}$ (by a vector space argument), and $\mathbb{Q}^{\kappa}$ is a homomorphic image of a product of finite cyclic groups (since $\mathbb{Q}$ is such an image). Thus $D$ is a homomorphic image of a product of finite cyclic groups, and as $R$ also satisfies this property, the group $\mathbb{A}=D\oplus R$ also has this property.
@@
We will first prove the equivalence of (1) and (2). Suppose that $A$ is cotorsion. By Proposition 10 we know that there is a sequence $(A_{i})_{i\in\omega}$ of abelian groups and a nonprincipal ultrafilter $\mathcal{U}$ on $\omega$ so that $A$ is a homomorphic image of the ultraproduct $(\prod_{i\in\omega}A_{i})/\mathcal{U}$. As $(\prod_{i\in\omega}A_{i})/\mathcal{U}$ is a homomorphic image of $(\prod_{i\in\omega}A_{i})/(\bigoplus_{i\in\omega}A_{i})$, so is $A$. For the other direction, if $A$ is a homomorphic image of a group of form $(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$ then $A$ is Higman-complete (Lemma 11), so $A$ is cotorsion (Proposition 10).
Now we argue that (1) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (1). That (1) implies (3) follows from the fact that a cotorsion group is a homomorphic image of an algebraically compact group (Proposition 10) and the fact that an algebraically compact abelian group is a homomorphic image of a product of finite cyclic groups (Lemma 16). That (3) implies (4) is evident. For the last implication, note that a finite group $F$ is Higman-complete. To see this, take a nonprincipal ultrafilter $\mathcal{V}$ on $\omega$ and recall that $F$ is isomorphic to the ultrapower $(\prod_{\omega}F)/\mathcal{V}$, so $F$ is Higman-complete (Proposition 11). Therefore a product of finite groups is Higman-complete (Lemma 13), and so $A$ is Higman-complete as the homomorphic image of a Higman-complete group (Remark 9). Thus $A$ is cotorsion (Proposition 10).
@@ 3. An analogue of a theorem of Chase
In the light of the very abelian nature of many of the previous results, we give a general method for producing (nonabelian) examples which give a negative solution to Question 3. These will be torsion groups which are not the homomorphic images of any nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{U}$. For this we will prove an analogue of a result of Chase [7].
@@
We’ll say a group $H$ is a bounded torsion group if there exists some $n\in\omega$ such that $(\forall h\in H)h^{n}=1_{H}$. For such an $H$ we let $\operatorname{ex}(H)\in\omega$ denote the smallest such $n$. If a torsion group is not bounded then we shall say it is an unbounded torsion group.
@@
Let $G$ be a topological group which is either
@@
and each $H_{j}$ a bounded torsion group and let $\phi:G\rightarrow\textbf{Bdd}\{H_{j}\}_{j\in J}$ be an abstract group homomorphism. Then there is some open neighborhood $U$ of the identity $1_{G}$ and $N\in\omega\setminus\{0\}$ such that $(\forall h\in\phi(U))h^{N}=1$.
@@
Let $H=\textbf{Bdd}\{H_{j}\}_{j\in J}$. Suppose by way of contradiction that the conclusion fails. The strategy will be to produce a system of equations (like in the definition of Higman-completeness) whose solution will give a contradiction. Suppose we are in situation (1) and let $d$ be a complete metric which induces the topology on $G$. We inductively choose
@@
For convenience, in the remainder of this section we’ll say a group which is a homomorphic image of a nonprincipal ultraproduct over $\omega$ is a ui-group (“ui” for ultraproduct image).
@@
If $\{H_{j}\}_{j\in J}$ is a collection of bounded torsion groups for which $\{\operatorname{ex}(H_{j})\}_{j\in J}$ is unbounded in $\omega$, then the direct sum $\bigoplus_{j\in J}H_{j}$ is not a ui-group.
Part of the attraction of Corollary 23 is that it provides many examples of non-ui-groups which are (unbounded) torsion. One can apply Corollary 23 to the collection $\{S_{j}\}_{j\in\omega}$ where $S_{j}$ is the symmetric group on $j$ elements, or to the collection $\{\mathfrak{b}(2,j)\}_{j\in\omega\setminus\{0,1\}}$ where $\mathfrak{b}(2,j)$ is the free Burnside group of rank $2$ and exponent $j$.
There exist ui-groups which are unbounded torsion. For example, a quasicyclic group $\mathbb{Z}(p^{\infty})$ is cotorsion and therefore a ui-group. We do not know the answers to the following.
@@
    If a ui-group is unbounded torsion then does it include a nontrivial divisible subgroup?
@@
    Is there a bounded torsion group which is not a ui-group?
@@
We give a few more examples of non-ui-groups. Recall that a group $H$ is cm-slender if every abstract group homomorphism from a completely metrizable group to $H$ has open kernel [6]. A cm-slender group is necessarily torsion-free, and examples include Baumslag-Solitar groups and nontrivial word-hyperbolic groups. Arguing as in Corollary 22 it is clear that a group having nontrivial homomorphic image in a cm-slender group is not a ui-group. Reasoning similarly, an infinite subgroup of the mapping class group of a connected compact surface is not a ui-group [4, Theorem 9.1], nor is a group which is nontrivially a free product of groups [13, Theorem 1.5].